The height of students in a class is distributed with mean and standard deviation. A random sample of 100 students was taken and the 90% confidence interval for mean was found to be between 175cm and 180cm.Estimate
i)value of the sample mean
Ii)value of standard deviation
Iii)95% confidence interval for mean
1
Expert's answer
2020-09-01T18:22:29-0400

The formula for confidence interval for the mean is given by,


"\bar x\ \pm\ Z(\alpha/2)*\frac {\sigma}{\sqrt N}"


Given that,

"\mu"1 = 180: Upper limit

"\mu"2 = 175: Lower limit

At 90% confidence interval "\alpha = 0.1"


"\mu"1 = 180 = "\bar x\ +\ Z(\alpha/2)*\frac {\sigma}{\sqrt N}................(Eq. \ 1)"


"\mu"2 = 175 = "\bar x\ -\ Z(\alpha/2)*\frac {\sigma}{\sqrt N}................(Eq.\ 2)"


N = 100


Hence,


"\mu1 - \mu2 = (\bar x+Z(\alpha/2)*\frac {\sigma}{\sqrt N}) - (\bar x-Z(\alpha/2)*\frac {\sigma}{\sqrt N})"


180-175 = "2*Z(\alpha/2)*\frac {\sigma}{\sqrt N}"


5 = 2 * Z0.05*"\frac{\sigma}{\sqrt 100}"





By using the standard normal distribution table we get the Z0.05 = 1.6449 and substituting it in the above equation we get,


"\sigma = 15.19849"


Substituting the value of mean in any one of the above equation (say eq. 1) we get,


180 = "\bar x +1.6449*\frac {15.19849}{\sqrt 100}"


Hence, "\bar x = 177.5"


ANSWER 1)


Sample mean "\bar x = 177.5"


ANSWER 2)


Standard deviation "\sigma = 15.19849"


ANSWER 3)


Using the above values for sample mean and standard deviation we can find 95% confidence interval for mean as under-


Here "\alpha = 0.05\ and\ (\alpha/2)\ = 0.025"





By using the standard normal distribution table we get the Z0.025 = 1.96 and substituting it in the above equation 1 & 2 we get,


"\mu1 =" 177.5 + 1.96*"\frac {15.19849}{\sqrt 100}" = 180.4789


"\mu2 =" 177.5 - 1.96*"\frac {15.19849}{\sqrt 100}" = 174.521


Hence the 95% confidence interval for mean is 174.521 cms & 180.4789 cms